# AuDeRe: Automated Strategy Decision and Realization in Robot Planning and Control via LLMs

### v1

@@ I Introduction
In this work, we propose a novel approach that leverages LLMs to intelligently select appropriate motion planning and control algorithms based on task-specific natural language instructions. Instead of directly generating trajectories or code, our framework uses LLMs to reason about task requirements, environmental constraints, and robot dynamics, subsequently deciding and invoking comprehensive planning and control application programming interfaces (APIs) tailored to these insights. We evaluate our method across diverse robotic scenarios with various complexity, ranging from simple trajectory tracking and basic planning tasks with collision avoidance to more complex scenarios involving maze navigation and high-level tasks encoded with signal temporal logic (STL) [20] specifications. Our approach is benchmarked against two baselines: an end-to-end LLM prediction method (LLM-predict) that directly outputs trajectories or control sequences, and a method (LLM-code) where the LLM generates executable code. Performance is quantitatively assessed using success rates, the average number of query iterations required for successful task completion, iteration-based success rates, and a detailed analysis of different error types. Our results show that the proposed LLM-driven approach (LLM-use-API) significantly outperforms these baselines in terms of higher success rates, fewer iterative queries, and reduced occurrence of errors across all tested scenarios.
@@ I-A Literature review
In [30], the authors introduced a framework employing few-shot prompts derived from physical environments, enabling LLMs to autoregressively predict low-level robotic control actions without necessitating task-specific fine-tuning. An interface presented in [13] seamlessly integrates natural language instructions with underlying model predictive control module. The work in [28] exploits prior knowledge of UAV dynamics encoded within an LLM, enabling dynamic adaptation of the entire control stack by adjusting low-level parameters, optimizing trajectory tracking commands, and supporting decision-making at the mission planning level. The approach outlined in [29] employs customizable few-shot ChatGPT prompts to transform natural language instructions and environmental information into executable, multi-stage robotic task plans supported by iterative feedback mechanisms. In [17], hierarchical prompting techniques are utilized to allow LLMs to autonomously generate robotic control code directly from natural language instructions. This structured approach effectively bridges semantic commands and executable robot actions. The integration presented in [1] pairs LLMs with pretrained robotic skill modules to ground semantic knowledge, enabling robots to interpret and execute intricate, temporally extensive natural language commands. The authors of [12] demonstrate how closed-loop linguistic feedback significantly enhances the reasoning and planning capabilities of LLMs within embodied robotic applications, eliminating the need for supplementary training. In [27], a structured, programmatic prompting method for robotic task planning is proposed, ensuring reliable generation of action sequences adaptable to specific environmental contexts. The methodology proposed in [8] combines LLMs with motion planning algorithms to achieve effective, human-aligned multi-object rearrangement tasks by leveraging commonsense reasoning. Text2Motion, as detailed in [18], is a language-based robotic planning system utilizing geometric feasibility heuristics, adept at solving complex sequential manipulation problems. The authors of [26] propose using natural language instructions to refine robot goals and constraints, demonstrating substantial improvements in planning effectiveness without real-time teleoperation. EUREKA, introduced in [19], is an LLM-driven method for designing reward functions tailored to complex manipulation tasks, significantly outperforming traditional human-crafted reward schemes across diverse reinforcement learning environments. The innovative framework proposed in [31] leverages LLMs to define effective reward structures, bridging the gap between high-level language instructions and low-level robotic control actions. In [10], natural language instructions are used for reward shaping within reinforcement learning, significantly enhancing sample efficiency in complex Atari environments. Finally, [4, 5] introduces an approach utilizing few-shot translation and autoregressive re-prompting in LLMs to translate natural language commands into intermediate representations, thus enabling robust integration with task and motion planning modules for advanced robotic operations.
@@ I-B Statement of contributions
In contrast to previous methods, our contributions are threefold. First, while lots of existing approaches either directly generate trajectories or rely solely on LLMs as code planners [7, 17], we utilize their advanced reasoning capabilities to intelligently select the most appropriate planning and control strategies by considering detailed task descriptions, environmental constraints, and robotic system dynamics. Second, our method innovatively leverages LLMs for automated and context-aware algorithm selection rather than limiting integration to a single predefined tool. To the best of our knowledge, this is the first work that allows LLMs to autonomously select strategies through available comprehensive planning and control APIs, rather than through a fixed integration framework. This novel approach enhances flexibility and adaptability to diverse task requirements. Finally, we introduce an iterative feedback mechanism through performance-based re-prompting, enabling continuous refinement of strategy decisions. To validate our approach, we conduct extensive experiments across scenarios of varying complexity, demonstrating significant performance improvements and generalizability. All together, these contributions yield a powerful and versatile framework that advances the state-of-the-art in automated robot planning and control.
The rest of the paper is organized as follows. In Section II, we formally describe the problem under consideration. Section III presents the detailed methods. The experimental results are provided in Section IV. Finally, Section V concludes the paper with remarks and suggestions for future work.
@@ II Problem statement
In this section, we introduce the overall problem setup, including system dynamics, objectives, and environmental constraints. Subsequently, we formulate the problem of leveraging an LLM-based automated strategy for robot planning and control.
@@
We define the robot’s workspace as $\mathcal{W}\subset\mathbb{R}^{n}$. Within this workspace, tasks are categorized into planning, control, or a combination of both. For control tasks, the reference trajectory $s_{r}$ is either defined as a set of waypoints in discrete time or as a mapping $s_{r}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n}$ for a continuous-time trajectory. Regarding planning tasks, obstacles are represented by the set $\mathcal{O}:=\{{\mathcal{O}_{1},\mathcal{O}_{2},\ldots,\mathcal{O}_{d}}\}$, with each obstacle $\mathcal{O}_{i}\subset\mathcal{W}$ indicating regions that must be avoided. Consequently, the collision-free workspace is given by $\mathcal{W}_{\text{free}}=\mathcal{W}\setminus\bigcup_{i=1}^{d}\mathcal{O}_{i}$. The primary objective is to navigate the robot from its initial state to the goal region $\mathcal{X}_{g}\subset\mathcal{W}_{\text{free}}$ while avoiding all obstacles within the workspace. In addition to spatial constraints, the natural language task $\mathcal{T}$ may also capture certain temporal constraints, which can be further represented as a high-level STL task. Hence, $\mathcal{T}$ generally ranges from simple tracking control or planning to complex planning and even high-level task specifications under spatiotemporal constraints. Rather than explicitly planning trajectories or manually selecting an established planning and control strategy, we leverage the reasoning capabilities of large language models to automatically determine an appropriate strategy. This decision-making process relies on problem descriptions expressed in natural language, which are inherently more intuitive and accessible for humans.
@@ III LLM-based strategy decision
In this section, we introduce an automated strategy decision process that leverages a large language model to select and iteratively refine planning and control strategies. First, we provide an overview of the approach, and then we detail each module.
@@ III-A Approach in a nutshell
The automated strategy decision process operates as follows and the illustration is shown in Fig. 2: Given a task $\mathcal{T}$ described in natural language, which encodes an environment setup comprising reference trajectories, initial conditions, goal regions, environmental constraints, the robot’s dynamics, and a set of available planning and control APIs denoted by $\mathcal{A}=\{\mathcal{A}_{1},\mathcal{A}_{2},\dots,\mathcal{A}_{N}\}$ is provided to the LLM. The LLM then selects an appropriate subset of these planning and control strategies, termed $\mathcal{A}^{\mathcal{T}}\subseteq\mathcal{A}$, to address the given task $\mathcal{T}$. Once the APIs are selected, the corresponding API code are provided back to the LLM. At this stage, the LLM generates a high-level execution plan, explicitly describing how the selected APIs will interact, detailing the input/output relations, and ensuring that data types and dimensions match across interfaces. The LLM configures the necessary parameters and interfaces, and generates the integration code required to execute the selected APIs. If errors occur, such as syntax errors or planning failures, an iterative refinement loop is triggered with a predefined maximum number of iterations. Through these iterative improvements, the final strategy and execution either complete the task or, if the maximum iterations are reached, report the errors.
@@ III-B Environment module
The environment module developed in this work provides a versatile and structured framework designed to facilitate the application and testing of LLMs for control and planning tasks. Specifically, it supports various dynamical systems, environmental constraints, initial state settings, and clearly defined target regions. The intended robot planning and control task, denoted by $\mathcal{T}$, is specified in natural language. The provided environmental functions and configurations comprehensively address the encoding and specification requirements associated with these natural-language-formulated problems.
The environment setup comprises multiple predefined dynamical models, including single and double integrators, unicycle, pendulum, and robotic-arm dynamics. These models allow comprehensive experiments across a range of planning and/or control scenarios. Each dynamical system is implemented across several computational frameworks such as PyTorch [21] for gradient-based control methods, and CasADi [3] for optimization and model predictive control methods, providing flexibility for integration with different planning and control APIs.
Environmental constraints are imposed through configurable obstacle types, including circles, squares, or mixed arrangements, while a boundary further defines the permissible workspace region $\mathcal{W}$. The module supports a broad range of task scenarios, spanning low-level trajectory tracking and basic state-space discretization to region graph navigation, planning within these graphs, and ultimately tracking generated reference trajectories. Initial states and goal regions can be randomized within structured constraints to ensure diverse path planning challenges. These scenarios with different complexities enable comprehensive evaluations of the robustness and generalization capabilities of LLM-driven planning and control policy selection, and can be applied to much more complex high-level planning tasks involving temporal logic specifications.
@@ III-C Planning-control API module
The planning-control module developed in this work provides a comprehensive framework designed to integrate seamlessly with LLM-driven methods for solving a wide range of planning and control tasks. The module offers a flexible and intuitive interface to address scenarios ranging from low-level trajectory tracking, state space discretization to high-level planning tasks involving complex temporal logic constraints.
To maximize versatility and effectiveness, the module provides eight distinct APIs, enabling an LLM to automatically select the most appropriate approach based on specific problem settings and constraints. The provided APIs and their functionalities are summarized below. They can be extended to include additional features, algorithms, and solvers:
@@
  • Gradient-based optimization (grad): Employs gradient descent and backpropagation using PyTorch for trajectory optimization, ensuring seamless integration of differentiable dynamics and loss functions.
@@
Each API is precisely defined with well-structured inputs and outputs, allowing for seamless integration and straightforward functionality for LLMs. Their diverse nature enables LLMs to flexibly and effectively address a wide range of planning-control challenges, providing strong generalization and adaptability across various task complexities.
@@ III-D LLM module and feedback loop
The LLM module serves as a crucial interface between high-level natural language descriptions of planning-control tasks and the detailed execution facilitated by the provided planning-control APIs. Initially, the input to the LLM consists of an environmental description, clearly defined task specifications, and a list of available planning-control APIs. Leveraging this context, the LLM decides the most appropriate APIs to fulfill the specified task.
@@
After the initial API selection, the LLM is prompted to retrieve detailed code for the corresponding APIs, including the required inputs, and expected outputs [16, 22]. Using this comprehensive information, it then generates executable code snippets that seamlessly interface with the selected APIs, configuring the necessary parameters and interfaces, and producing the integration code required for execution.
The refinement process involves iterative feedback loops to ensure correctness, robustness, and optimal performance. Generated code undergoes syntax checking and planner checks to verify compliance with specified goals, environmental constraints, and trajectories for tracking. Any spotted errors or suboptimal implementations trigger further iterative refinements by the LLM, in order to progressively improving the code quality and functional correctness [4, 23].
Through these iterative refinement cycles, which incorporate syntax validation and trajectory planning assessments, the final LLM-based strategy converges to reliable planning and/or closed-loop trajectories that accurately and effectively fulfill the defined planning and control tasks. The entire LLM-based strategy decision process and the roles of LLMs are illustrated by Fig. 2.
Based on the above discussions for each module, we present an example on LLM prompt in Fig. 3 for a simple tracking control problem, where the symbols * or ***∗ ∗ ∗ denote detailed parameters and text that can be omitted.
@@ IV Experiments
In this section, we present the experimental setup designed to evaluate the performance of our proposed method. We first describe the specific scenarios considered, outline the baseline methods selected for comparison, and detail the metrics used to assess the performance. Finally, we demonstrate through extensive experiments results that our method consistently outperforms these baselines across various scenarios.
@@ IV-A Scenarios
    Complex planning in a 3 by 3 maze with extensive collision environment: A more intricate planning scenario where the robot must efficiently navigate through a structured 3 by 3 maze while avoiding multiple obstacles distributed throughout the environment.
@@
For generalizability, each scenario is extensively evaluated over $100$ experiments. We randomize environmental parameters such as reference trajectories, initial states, goal regions, and obstacle placements, ensuring diverse and representative conditions for a thorough comparison. Fig. 4 shows one successful experiment per scenario. In the simple tracking problems (a) and (b), the “mpc” API is called. For the simple planning scenario (c), the “rrt” API is employed. In the complex 3 by 3 maze planning case, a combination of the “astar” and “rrt” APIs is applied. For the high-level STL task in (e), the “milp” API is invoked. All experiments are implemented using the GPT-4o large language model.

### v2

@@ I Introduction
Traditional planning and control algorithms provide strong theoretical guarantees when carefully chosen and tuned for a specific task. However, in complex and dynamic environments, selecting suitable strategies is a nontrivial challenge that requires significant expertise and fails to adapt to changes or scale effectively. To address this limitation, we propose a novel approach that leverages LLMs to select motion planning and control algorithms based on task descriptions in order to reduce human effort, improve adaptability, and enable broader applicability across diverse tasks. Instead of directly generating trajectories or code, our framework uses LLMs to reason about task requirements, environmental constraints, and robot dynamics, subsequently deciding and invoking comprehensive planning and control application programming interfaces (APIs) tailored to these insights. We evaluate our method across diverse robotic scenarios with various complexity, ranging from simple trajectory tracking and basic planning tasks with collision avoidance to more complex scenarios involving maze navigation and high-level tasks encoded with signal temporal logic (STL) [19] specifications. Our approach is benchmarked against two baselines: an end-to-end LLM prediction method (LLM-predict) that directly outputs trajectories or control sequences, and a method (LLM-code) where the LLM generates executable code. Performance is quantitatively assessed using success rates, the average number of query iterations required for successful task completion, iteration-based success rates, and a detailed analysis of different error types. Our results show that the proposed LLM-driven approach (LLM-use-API) significantly outperforms these baselines in terms of higher success rates, fewer iterative queries, and reduced occurrence of errors across all tested scenarios.
@@ I-A Literature review
In [29], the authors introduce a framework that uses few-shot prompts derived from physical environments. This framework allows LLMs to autoregressively predict low-level robotic control actions without requiring task-specific fine-tuning. An interface presented in [12] integrates natural language instructions with underlying model predictive control module. The work in [27] exploits prior knowledge of UAV dynamics encoded within an LLM. It enables dynamic adaptation of the entire control stack by adjusting low-level parameters, optimizing trajectory tracking commands, and supporting decision-making at the mission planning level. The approach outlined in [28] employs customizable few-shot prompts to transform natural language instructions and environmental information into executable, multi-stage robotic task plans supported by iterative feedback mechanisms. In [16], hierarchical prompting techniques are utilized to allow LLMs to autonomously generate robotic control code directly from natural language instructions. This structured approach effectively bridges semantic commands and executable robot actions. The integration in [1] combines LLMs with pretrained robotic skill modules to ground semantic knowledge, which allows robots to interpret and execute complex natural language commands. The authors of [11] demonstrate how closed-loop linguistic feedback significantly enhances the reasoning and planning capabilities of LLMs within embodied robotic applications. In [26], the authors propose a structured programmatic prompting method for robotic task planning that ensures reliable generation of action sequences adapted to specific environmental contexts. The methodology proposed in [7] combines LLMs with motion planning algorithms to achieve effective human-aligned multi-object rearrangement tasks utilizing common sense reasoning. Text2Motion, as described in [17], is a language-based robotic planning system that employs geometric feasibility heuristics and effectively solves complex sequential manipulation problems. The authors of [25] propose the use of natural language instructions to refine robot goals and constraints, and show substantial improvements in planning effectiveness without reliance on real-time teleoperation. EUREKA, introduced in [18], is an LLM-driven method for the design of reward functions tailored to complex manipulation tasks, and it clearly outperforms traditional human-crafted reward schemes across various reinforcement learning environments. The framework in [30] leverages LLMs to define effective reward structures and creates a connection between high-level language instructions and low-level robotic control actions. In [9], natural language instructions serve as a tool for reward shaping in reinforcement learning, and this approach leads to benefits in sample efficiency for complex Atari environments. Finally, [4] introduces an approach based on few-shot translation and autoregressive re-prompting in LLMs to translate natural language commands into intermediate representations, which enables robust integration with task and motion planning modules for robotic operations.
@@ I-B Statement of contributions
In contrast to previous methods, our contributions are threefold. First, while lots of existing approaches either directly generate trajectories or rely solely on LLMs as code planners [6, 16], we utilize their advanced reasoning capabilities to intelligently select the appropriate planning and control strategies by considering detailed task descriptions, environmental constraints, and robotic system dynamics. Second, our method innovatively leverages LLMs for automated and context-aware algorithm selection rather than limiting integration to a single predefined tool. To the best of our knowledge, this is the first work to frame LLM-based robot planning and control as an automated strategy decision problem across multiple comprehensive planning and control APIs, rather than as direct trajectory generation or integration with a single fixed tool. This novel approach enhances flexibility and adaptability to diverse task requirements. Finally, we introduce an iterative feedback mechanism through performance-based re-prompting, enabling continuous refinement of strategy decisions. To validate our approach, we conduct extensive experiments across scenarios of varying complexity, demonstrating significant performance improvements and generalizability. All together, these contributions yield a powerful and versatile framework that advances the state-of-the-art in automated robot planning and control.
@@ II Problem statement
In this section, we introduce the problem setup, including system dynamics, objectives, and environmental constraints. Then, we formulate the problem of leveraging an LLM-based automated strategy for robot planning and control.
@@
We define the robot’s workspace as $\mathcal{W}\subset\mathbb{R}^{n}$, and the tasks are categorized into planning, control, or a combination of both. For control tasks, the reference trajectory $s_{r}$ is either defined as a set of waypoints in discrete time or as a mapping $s_{r}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n}$ for a continuous-time trajectory. Regarding planning tasks, obstacles are represented by the set $\mathcal{O}:=\{{\mathcal{O}_{1},\mathcal{O}_{2},\ldots,\mathcal{O}_{d}}\}$, with each obstacle $\mathcal{O}_{i}\subset\mathcal{W}$ indicating regions that must be avoided. The collision-free workspace is therefore given by $\mathcal{W}_{\text{free}}=\mathcal{W}\setminus\bigcup_{i=1}^{d}\mathcal{O}_{i}$. The main objective is to navigate the robot from its initial state to the goal region $\mathcal{X}_{g}\subset\mathcal{W}_{\text{free}}$ while avoiding all obstacles. In addition to spatial constraints, the natural language task $\mathcal{T}$ may also capture certain temporal constraints, which can be further represented as a high-level STL task. Hence, $\mathcal{T}$ generally ranges from simple tracking control or planning to complex planning and even high-level task specifications under spatiotemporal constraints. Rather than explicitly planning trajectories or manually selecting an established planning and control strategy, we leverage the reasoning capabilities of large language models to automatically determine an appropriate strategy. This decision-making process relies on problem descriptions expressed in natural language, which are inherently more intuitive and accessible for humans.
@@ III LLM-based strategy decision
In this section, we introduce an automated strategy decision process that leverages an LLM to select and iteratively refine planning and control strategies. First, we provide an overview of the approach, and then we detail each module.
@@ III-A Approach in a nutshell
The automated strategy decision process operates as follows and the illustration is shown in Fig. 2: Given a task $\mathcal{T}$ described in natural language, which encodes an environment setup comprising reference trajectories, initial conditions, goal regions, environmental constraints, the robot’s dynamics, and a set of available planning and control APIs denoted by $\mathcal{A}=\{\mathcal{A}_{1},\mathcal{A}_{2},\dots,\mathcal{A}_{N}\}$ is provided to the LLM. The LLM then selects an appropriate subset of these planning and control strategies, termed $\mathcal{A}^{\mathcal{T}}\subseteq\mathcal{A}$, to address the given task $\mathcal{T}$. Once the APIs are selected, the corresponding API codes are provided back to the LLM. At this stage, the LLM generates a high-level execution plan, explicitly describing how the selected APIs will interact, detailing the input/output relations, and ensuring that data types and dimensions match across interfaces. The LLM configures the necessary parameters and interfaces, and generates the integration code required to execute the selected APIs. If errors occur, such as syntax errors or planning failures, an iterative refinement loop is triggered with a predefined maximum number of iterations. Through these iterative improvements, the final strategy and execution either complete the task or, if the maximum iterations are reached, report the errors.
@@ III-B Environment module
The environment module developed in this work provides a versatile and structured framework designed to facilitate the application and testing of LLMs for control and planning tasks. Specifically, it supports various dynamical systems, environmental constraints, initial states, and clearly defined target regions. The intended robot planning and control task, denoted by $\mathcal{T}$, is specified in natural language. The provided environmental functions and configurations comprehensively address the encoding and specification requirements associated with these problems formulated in natural language.
The environment setup comprises multiple predefined dynamical models, including single and double integrators, unicycle, pendulum, and robotic-arm dynamics. These models allow comprehensive experiments across a range of planning and control scenarios. Each dynamical system is implemented across several computational frameworks such as PyTorch [20] for gradient-based control methods, and CasADi [3] for optimization and model predictive control methods, providing flexibility for integration with different planning and control APIs.
Environmental constraints are imposed through configurable obstacle types, including circles, squares, or mixed arrangements, while a boundary further defines the permissible workspace region $\mathcal{W}$. The module supports a broad range of task scenarios, spanning low-level trajectory tracking and basic state space discretization to region graph navigation, planning within these graphs, and ultimately tracking generated reference trajectories. Initial states and goal regions can be randomized within structured constraints to ensure diverse path planning challenges. These scenarios with different complexities enable comprehensive evaluations of the robustness and generalization capabilities of LLM-driven planning and control policy selection, and can be applied to much more complex high-level planning tasks involving temporal logic specifications.
@@ III-C Planning-control API module
The planning-control module developed in this work provides a comprehensive framework designed to integrate effectively with LLM-driven methods for a wide range of planning and control tasks. The module offers a flexible and intuitive interface to address scenarios ranging from low-level trajectory tracking, state space discretization to high-level planning involving complex temporal logic constraints.
To maximize versatility and effectiveness, the module provides eight distinct APIs, enabling an LLM to automatically select the appropriate approach based on specific problem settings and constraints. The provided APIs and their functionalities are summarized below. Our API set and task environments are chosen for clarity of evaluation, but the framework is not limited to these implementations. The APIs are modular and can be extended with additional planners, controllers, solvers, or external libraries. The LLM reasoning layer is environment-agnostic, so new algorithms can be added by exposing the relevant APIs. This design emphasizes the framework’s generalizability beyond the presented scenarios, as demonstrated in Sec. IV-A.
@@
  • Gradient-based optimization (grad): Employs gradient descent and backpropagation using PyTorch for trajectory optimization, ensuring efficient integration of differentiable dynamics and loss functions.
@@
Each API is precisely defined with well-structured inputs and outputs, which allows efficient integration and straightforward functionality for LLMs. Their diverse nature enables LLMs to flexibly and effectively address a wide range of planning-control challenges. This provides strong generalization and adaptability across tasks of varying complexity.
@@ III-D LLM module and feedback loop
The LLM module serves as a crucial interface between high-level natural language task descriptions and the detailed execution facilitated by the provided APIs. Initially, the input to the LLM consists of an environmental description, clearly defined task specifications, and a list of available planning-control APIs. Leveraging this context, the LLM decides the appropriate APIs to fulfill the specified task.
After the initial API selection, the LLM is prompted to retrieve detailed code for the corresponding APIs, including the required inputs, and expected outputs [15, 21]. Using this comprehensive information, it then generates executable code snippets that interface with the selected APIs, configuring the necessary parameters and interfaces, and producing the integration code required for execution.
@@
The refinement process involves iterative feedback loops to ensure correctness and improved performance. Generated code undergoes syntax checking and planner checks to verify compliance with specified goals, environmental constraints, and trajectories for tracking. Any spotted errors or implementations that fail to meet requirements trigger further iterative refinements by the LLM, in order to progressively improve the code quality and functional correctness [4, 22]. We present an example on LLM prompt in Fig. 3 for a simple tracking control problem, where the symbols * or *** denote detailed parameters and text that can be omitted. In this example, we include the diagnostic summary from the previous round in the new prompt, and let LLM refine its strategy based on past performance feedback.
Through these iterative refinement cycles, which incorporate syntax validation and trajectory planning assessments, the final LLM-based strategy converges to reliable planning and closed-loop trajectories that accurately and effectively fulfill the tasks. The entire LLM-based strategy decision process and the roles of LLMs are illustrated by Fig. 2.
@@ IV Experiments
In this section, we present the experimental setup designed to evaluate the performance of our proposed method. We first describe the scenarios considered, outline the baseline methods selected for comparison, and detail the metrics used to assess the performance. Finally, we demonstrate through extensive experiments results that our method consistently outperforms these baselines across various scenarios.
@@ IV-A Scenarios
    Complex planning in a $3\times 3$ maze with extensive collision environment: A more intricate planning scenario where the robot must efficiently navigate through a structured $3\times 3$ maze while avoiding multiple obstacles distributed throughout the environment.
@@
For generalizability, each scenario is extensively evaluated over $100$ experiments. We randomize environmental parameters such as reference trajectories, initial states, goal regions, and obstacle placements, ensuring diverse and representative conditions for a thorough comparison. Fig. 4 shows one successful experiment per scenario. In the simple tracking problems (a) and (b), the “mpc” API is called. For the simple planning scenario (c), the “rrt” API is employed. In the complex $3\times 3$ maze planning case, a combination of the “astar” and “rrt” APIs is applied. For the high-level STL task in (e), the “milp” API is invoked. All experiments are implemented using the GPT-4o large language model.